Suppose that the probability that a corn seed from a certain batch does not germinate equals 0.02. If we plant 200 of these seeds, what is the probability that

• A) At most 5 seeds will not germinate?

• B) Exactly 3 will not germinate?

• C) At least 3 will not germinate?



1
Expert's answer
2022-06-01T13:08:25-0400

"X\sim Bin(n=200,p=0.02)"

Check if"X" has approximately a normal distribution with "\mu=np" and "\sigma=\sqrt{npq}." .

In practice, the approximation is adequate provided that both "np\ge10" and "nq\ge10," since there is then enough symmetry in the underlying binomial

distribution.

Check


"np=200(0.02)=4<10"

When the value of "n"  in a binomial distribution is large and the value of "p"  is very small, the binomial distribution can be approximated by a Poisson distribution. If "n > 20" and "np < 5" or "nq < 5" then the Poisson is a good approximation.

Check


"n=200>20, np=200(0.02)=4<10"


Then "\lambda=np=4"

"X\sim Po(4)"


A)


"P(X\le5)=P(X=0)+P(X=1)"

"+P(X=2)+P(X=3)"

"+P(X=4)+P(X=5)"

"=\dfrac{e^{-4}(4)^0}{0!}+\dfrac{e^{-4}(4)^1}{1!}+\dfrac{e^{-4}(4)^2}{2!}"

"+\dfrac{e^{-4}(4)^3}{3!}+\dfrac{e^{-4}(4)^4}{4!}+\dfrac{e^{-4}(4)^5}{5!}"

"=0.78513"

B)


"P(X=3)=\dfrac{e^{-4}(4)^3}{3!}=0.19537"

C)


"P(X\ge3)=1-P(X=0)-P(X=1)"

"-P(X=2)=1-\dfrac{e^{-4}(4)^0}{0!}-\dfrac{e^{-4}(4)^1}{1!}"

"-\dfrac{e^{-4}(4)^2}{2!}=0.76190"